IMPORTANT Questions 1 & 2 ask for cash flows only, no present values. They are a critical part of the problem, but since the116 Mon 2: Tixta Smrt Prom Table 1: Present Value of $1.00 perind 1129 13% 14% 5% 6% 7% 0.980 1 0.971 0.990 06 10052 0.962 0.
Answer #1

cash flow for current system

Total defective units-detected

=2000-1600

=400units

year 3 cash flow

Annual inspection cost -repairing cost *units repaired- refunded units

= -40000 -(75*1600)-400*195

=-40000-120000-78000

= -238000$

new system cash flow in year 3

annual inspection cost -repairing cost*units repaired - refunded units

= -[40000+21000increase]-330*48-[400-330]*195*120%

= -61000-15840-16380

= -93220

NPV

CURRENT SYSTEM

YEAR cash flow pv factor at 6% presnt value
1-4 -238000 3.465 -824670$ [238000*3.465]

new system

YEAR cash flow pv factor at 6% presnt value
1-4 -93220 3.465 -323007
0 -190000 1 -190000
0 -470000 1 -470000
4 21000 0.792 16632
NET PRESENT VALUE -966375$